OpenStudy (anonymous):

If r = 4 units and x = 16 units, then what is the volume of the cylinder shown above? 256 cubic units 64 cubic units 1,024 cubic units 160 cubic units

OpenStudy (anonymous):

Where's the picture?

OpenStudy (anonymous):

Save it to your computer and upload it here using the 'Attach File' button.

OpenStudy (anonymous):

Lol, that works

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Okay, so the formula is: \(\sf V = \pi r^2 h\) Plug in what we know: \(\sf V = (3.14)(4^2)(16)\) Can you multiply those?

OpenStudy (anonymous):

\(\sf 4^2 \rightarrow 4 \times 4 \rightarrow 16\) So we have: \(\sf V = (3.14)(16)(16)\) Can you multiply 3.14 * 16 * 16? @Icecube75789

OpenStudy (anonymous):

803.84

OpenStudy (anonymous):

Yes, but I think it wants the answer in terms of pi.. \(\sf V = (\pi)(16)(16)\) So what's 16 * 16?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Yep, your answer is \(\sf 256 \pi\).

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @Icecube75789

OpenStudy (anonymous):

Ok thank u